OpenStudy (anonymous):

Find the slope of the line that passes through each pair of points. (-3,-3),(-7,-5)

OpenStudy (anonymous):

m = rise/run -5 - (-3) = -2 -7 - (-3) = -4 -2/-4 = 1/2 m = 1/2

OpenStudy (anonymous):

Can you please explain I do not understand.

OpenStudy (lgbasallote):

slope = \[\frac{y_2 - y_1}{x_2 - x_1}\] y2 means y coordinate of the second ordered pair y1 is the first y x1 is the first x x2 is the second x therefore \[slope = \frac{-5 - (-3)}{-7 - (-3)}\] \[slope = \frac{-5 + 3}{-7 + 3}\] \[slope = \frac{-2}{-4}\] \[slope = \frac{1}{2}\]

OpenStudy (lgbasallote):

got it @rosimary ?

OpenStudy (anonymous):

Hi Sorry. Ok so to make sure I understand. Because I wrote the points wrong. (-9,-3),(-7,-5) m= -2/2 wich then turns to -1 ? Yes?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Great. Thank you so much. I do have other questions. could you help?

OpenStudy (lgbasallote):

sadly i have a class in 12 minutes so im out in 2 minutes sorry

OpenStudy (anonymous):

ok thanks